# P-MapNet: Far-seeing Map Generator Enhanced by both SDMap and HDMap Priors

### v2

@@ Abstract
Autonomous vehicles are gradually entering city roads today, with the help of high-definition maps (HDMaps). However, the reliance on HDMaps prevents autonomous vehicles from stepping into regions without this expensive digital infrastructure. This fact drives many researchers to study online HDMap generation algorithms, but the performance of these algorithms at far regions is still unsatisfying. We present P-MapNet, in which the letter P highlights the fact that we focus on incorporating map priors to improve model performance. Specifically, we exploit priors in both SDMap and HDMap. On one hand, we extract weakly aligned SDMap from OpenStreetMap, and encode it as an additional conditioning branch. Despite the misalignment challenge, our attention-based architecture adaptively attends to relevant SDMap skeletons and significantly improves performance. On the other hand, we exploit a masked autoencoder to capture the prior distribution of HDMap, which can serve as a refinement module to mitigate occlusions and artifacts. We benchmark on the nuScenes and Argoverse2 datasets. Through comprehensive experiments, we show that: (1) our SDMap prior can improve online map generation performance, using both rasterized (by up to $+18.73$ $\rm mIoU$) and vectorized (by up to $+8.50$ $\rm mAP$) output representations. (2) our HDMap prior can improve map perceptual metrics by up to $6.34\%$. (3) P-MapNet can be switched into different inference modes that covers different regions of the accuracy-efficiency trade-off landscape. (4) P-MapNet is a far-seeing solution that brings larger improvements on longer ranges. Codes and models are publicly available at https://jike5.github.io/P-MapNet/.
@@ II-B Long-range Map Perception
To enhance the practicality of HD maps for downstream tasks, some studies aim to extend their coverage to longer perception ranges. SuperFusion[5] combines LiDAR point clouds and camera images for depth-aware BEV transformation, yielding forward-view HD map predictions up to $90~{}m$. NeuralMapPrior[27] maintains and updates a global neural map prior, enhancing online observations to generate higher-quality, extended-range HD map predictions. [8] proposes using satellite maps to aid online map generation. Features from onboard sensors and satellite images are aggregated through a hierarchical fusion module to obtain the final BEV features. MV-Map[26] specializes in offline, long-range HD map generation. It aggregates all relevant frames during traversal and optimizes neural radiance fields for improved BEV feature generation.
@@ III Formulation
$\mathcal{M}=\mathcal{F}_{2}(\mathcal{F}_{1}(\mathcal{P},\mathcal{I})),\$ (1)
@@
$\mathcal{M}_{S}=\mathcal{F}_{2}(\mathcal{F}_{1}(\mathcal{P},\mathcal{I},\mathcal{S})),\$ (3)
@@ IV-B HDMap Prior Module
Figure 4: Qualitative results. We conduct a comparative analysis within a range of 240m×$\times$60m on nuScenes dataset and 120m×$\times$60m on Argoverse2 dataset, utilizing C+L as input. In our notation, “S” indicates that our method utilizes only the SDMap priors, while “S+H” indicates the utilization of both. Our method consistently outperforms the baseline method under various weather conditions and in scenarios involving viewpoint occlusion.
@@ A-B1 Integrating SDMap Prior into end-to-end vectorized framework
The visualization results in Fig. 6 also show that MapTR-SDMap performs better under the most challenging $240\times 60m$ ultra-long range of perception. It can also be seen that the segmentation-post-processing approach has stable results because it is sense-prediction, while the end-to-end vectorization approach still has some challenges such as significant predictive bias and challenges in keypoint selection. In conclusion, our proposed SDMap Prior fusion method demonstrates performance improvement in both the segmentation-postprocessing framework and the end-to-end framework.
Figure 6: The qualitative results of vectorized results on a perception range of $240m\times 60m$ We integrate the SDMap Prior Module into the MapTR (with only minor modifications), referred to as the MapTR-SD. PMapNet-CAM is our method with both SDMap prior and HDMap prior utilizing the post process.

### v3

@@ Abstract
Autonomous vehicles are gradually entering city roads today, with the help of high-definition maps (HDMaps). However, the reliance on HDMaps prevents autonomous vehicles from stepping into regions without this expensive digital infrastructure. This fact drives many researchers to study online HDMap generation algorithms, but the performance of these algorithms at far regions is still unsatisfying. We present P-MapNet, in which the letter P highlights the fact that we focus on incorporating map priors to improve model performance. Specifically, we exploit priors in both SDMap and HDMap. On one hand, we extract weakly aligned SDMap from OpenStreetMap, and encode it as an additional conditioning branch. Despite the misalignment challenge, our attention-based architecture adaptively attends to relevant SDMap skeletons and significantly improves performance. On the other hand, we exploit a masked autoencoder to capture the prior distribution of HDMap, which can serve as a refinement module to mitigate occlusions and artifacts. We benchmark on the nuScenes and Argoverse2 datasets. Through comprehensive experiments, we show that: (1) our SDMap prior can improve online map generation performance, using both rasterized (by up to +18.73+18.73 $\rm mIoU$) and vectorized (by up to +8.50+8.50 $\rm mAP$) output representations. (2) our HDMap prior can improve map perceptual metrics by up to $6.34\%$. (3) P-MapNet can be switched into different inference modes that covers different regions of the accuracy-efficiency trade-off landscape. (4) P-MapNet is a far-seeing solution that brings larger improvements on longer ranges. Codes and models are publicly available at https://jike5.github.io/P-MapNet/.
@@ II-B Long-range Map Perception
To enhance the practicality of HD maps for downstream tasks, some studies aim to extend their coverage to longer perception ranges. SuperFusion[5] combines LiDAR point clouds and camera images for depth-aware BEV transformation, yielding forward-view HD map predictions up to 90m90~m. NeuralMapPrior[27] maintains and updates a global neural map prior, enhancing online observations to generate higher-quality, extended-range HD map predictions. [8] proposes using satellite maps to aid online map generation. Features from onboard sensors and satellite images are aggregated through a hierarchical fusion module to obtain the final BEV features. MV-Map[26] specializes in offline, long-range HD map generation. It aggregates all relevant frames during traversal and optimizes neural radiance fields for improved BEV feature generation.
@@ III Formulation
$\mathcal{M}=\mathcal{F}_{2}(\mathcal{F}_{1}(\mathcal{P},\mathcal{I})),\\$ (1)
@@
$\mathcal{M}_{S}=\mathcal{F}_{2}(\mathcal{F}_{1}(\mathcal{P},\mathcal{I},\mathcal{S})),\\$ (3)
@@ IV-B HDMap Prior Module
[…]
@@ A-B1 Integrating SDMap Prior into end-to-end vectorized framework
The visualization results in Fig. 6 also show that MapTR-SDMap performs better under the most challenging $240\times 60m$ ultra-long range of perception. It can also be seen that the segmentation-post-processing approach has stable results because it is dense-prediction, while the end-to-end vectorization approach still has some challenges such as significant predictive bias and challenges in keypoint selection. In conclusion, our proposed SDMap Prior fusion method demonstrates performance improvement in both the segmentation-postprocessing framework and the end-to-end framework.
Figure 6: The qualitative results of vectorized results on a perception range of $240m\times 60m$ . We integrate the SDMap Prior Module into the MapTR (with only minor modifications), referred to as the MapTR-SD. PMapNet-CAM is our method with both SDMap prior and HDMap prior utilizing the post process.